# Learning to Steer Learners in Games

### v1

@@ 6.2 Learning to Steer Mirror Ascent Learners
$y_{t+1}=\arg\min_{y\in\Delta_{n}}\left\{-\eta_{t}D(y\|y_{t})-(x_{t}^{T}B+\xi_{t}^{T})y\right\}$ (25)
where $\xi_{t}\in\mathbb{R}^{n}$ is some noise that is either innate in the problem or injected by $P_{2}$ to prevent from information leakage. We assume that $\eta_{t}$ and $D(\cdot\|\cdot)$ are both known to $P_{1}$ and the regularizer satisfies $\nabla_{y}D(y_{t+1}\|y_{t})\rightarrow\infty$ if there exists $i\in[n]$ such that $y_{t+1,i}\rightarrow 0$.
@@ B.1 Proof of Proposition 1
B.2 Proof of Theorem 2
@@
where we have used the fact that $1-y_{t,i}=\sum_{j\in[n],j\neq i}y_{t,j}$ for all $y_{t}\in\Delta_{n}$ in the fourth equation. Since $\mathcal{A}_{2}$ used by the learner is $f$-no-regret, we have:
@@
$=$ $O(f(T)/d_{2}),$
where the second equality follows from $y_{t,i}\in[0,1],\forall y_{t}\in\Delta_{n},i\in[n]$ and the third inequality follows from the same argument as above. Let $i^{*}$ denote $\operatorname*{arg\,max}_{i}V_{i}(A,B)$, the Stackelberg regret of $P_{1}$ satisfies:
@@
$=$ $\left(V_{i^{*}}(A,B)-V_{i^{*}}^{-}(A,B)\right)+\|A^{T}x^{-}\|_{\infty}\|\sum_{t=1}^{T}(e_{i^{*}}-y_{t})\|_{1}$
$\stackrel{{\scriptstyle\text{(iii)}}}{{\leq}}$ $Td_{1}\|A\|_{\max}+O(f(T)/d_{2}),$
where we have used Hölder’s inequality in (i), the maximizing argument of (10) in (ii) and Proposition 1 in (iii).
@@ C.4 Proof of Lemma 3
C.5 Proof of Theorem 4
@@
$\text{maximize}_{i,x}$ $V^{-}_{i}(A,\hat{\mathcal{B}})=x^{T}a_{i}$ (80)
@@
$\leq$ $4(g(T)+\sqrt{Tf(T)})\|A_{:,i^{*}}\|_{\infty}Sen((\mathcal{B}_{i^{*}}^{\circ})^{T})+\|A\|_{\max}\frac{C\sqrt{Tf(T)}}{\max_{j,k}\|B(e_{j}-e_{k})\|_{\infty}}$
@@ Appendix E Algorithms and Proofs for Section 6
Let $\tilde{y}\in BR(B,\tilde{x})$ be a best response to $\tilde{x}$. For the ”if” direction, on observing that:
@@
for some constant $C^{\prime}$ where we have used the fact that $\tau=O(f(T))$, and therefore $\{x_{t},y_{t}\}_{t=1}^{T}$ is $f$-no-regret. For the ”only if” direction, notice that
@@
E.1 Algorithm 1, Proof of Theorem 5 and Discussion
@@
Proof of Theorem 5.
@@
Now we analyze the Stackelberg regret against an $f$-no-regret algorithm $\mathcal{A}_{2}$. W.l.o.g suppose $e_{2}$ is the strictly dominated action, the assumption that each facet has length at least $d$ indicates that there exists a constant $\epsilon$, such that:
$\tilde{x}Be_{1}-\tilde{x}Be_{2}\geq\epsilon.$ (103)
@@
$\|\sum_{t=5}^{\infty}(e_{1}-y_{t})\|_{\infty}\leq\frac{C}{\epsilon}f(T).$ (105)
@@
$StackReg_{1}(\mathcal{A}_{1},\mathcal{A}_{2})$ $\leq 4V(A,B)-\sum_{t=1}^{4}x_{t}^{T}Ay_{t}+\tilde{x}^{T}B\sum_{t=5}^{T}(e_{1}-y_{t})$ (106)
$\leq 4\|A\|_{\max}+\|\tilde{x}^{T}B\|_{1}\|\sum_{t=5}^{T}(e_{1}-y_{t})\|_{\infty}$
$\leq 4\|A\|_{\max}+\|B\|_{\max}\frac{C}{\epsilon}f(T)$
$=O(f(T)).$
If neither facets are empty, the binary search phase shrinks the interval length from $1$ to $d$, which requires at most $-\log d+1$ calls of the BinarySearch function. Each Binary search calls the procedure test for at most one time, which corresponds to at most two interaction time steps. Therefore, the total number of time steps in the binary search phase is at most $-2\log d+C$ for some constant $C$, leading to a Stackelberg regret of at most $-2\log d+C$, which is $O(f(T))$.
@@
$StackReg_{1}(\mathcal{A}_{1},\mathcal{A}_{2})$ $\leq\lfloor-2\log d+C\rfloor V(A,B)-\sum_{t=1}^{\lfloor-2\log d+C\rfloor}x_{t}^{T}Ay_{t}+\tilde{x}^{T}B\sum_{t=\lfloor-2\log d+C\rfloor+1}^{T}(e_{1}-y_{t})$ (107)
$\leq(-2\log d+C)\|A\|_{\max}+\|\tilde{x}^{T}B\|_{1}\|\sum_{t=-2\log d+C}^{T}(e_{1}-y_{t})\|_{\infty}$
$\leq(-2\log d+C)\|A\|_{\max}+O(\frac{f(T)}{d}+dT)$
$=O(\frac{f(T)}{d}+dT-\log d).$
@@
E.2 Algorithm 4 and Proof of Theorem 6
@@
Proof of Theorem 6.
@@
$L(y,\lambda,\mu)=-\eta_{t}D(y\|y_{t})-(x_{t}^{T}B+\xi_{t}^{T})y-\lambda^{T}y+\mu(\sum_{i}y_{i}-1).$ (113)
The KKT condition yields:
$-\eta_{t}\nabla_{t+1}yD(y_{t+1}\|y_{t})-B^{T}x_{t}-\lambda+\mu\mathrm{1}-\xi_{t}=0;$ (114)
@@
Since the regularizer satisfies $\nabla_{y}D(y_{t+1}\|y_{t})\rightarrow\infty$ if there exists $i\in[n]$ such that $y_{t+1,i}\rightarrow 0$, we can deduce that $\lambda=0$, and the condition becomes:
$-\eta_{t}\nabla_{y}D(y_{t+1}\|y_{t})-B^{T}x_{t}+\mu\mathrm{1}-\xi_{t}=0.$ (115)
Let $h_{t}$ denote $h_{t}:=-\eta_{t}\nabla_{y}D(y_{t+1}\|y_{t})$, we know that
@@
$StackReg_{1}(\mathcal{A}_{1},\mathcal{A}_{2})$ $\leq m(k+1)V(A,B)+O(\sqrt{Tf(T)}+g(T))$ (126)
@@
Appendix F Proof of Auxiliary Lemmas
F.1 Proof of Lemma 7

### v2

@@ 6.2 Learning to Steer Mirror Ascent Learners
$y_{t+1}=\arg\min_{y\in\Delta_{n}}\left\{\eta_{t}D(y\|y_{t})-(x_{t}^{T}B+\xi_{t}^{T})y\right\}$ (26)
where $\xi_{t}\in\mathbb{R}^{n}$ is some noise that is either innate in the problem or injected by $P_{2}$ to prevent from information leakage. We assume that $\eta_{t}$ and the Bregman divergence regularizer $D(\cdot\|\cdot)$ are both known to $P_{1}$ and the regularizer satisfies $\nabla_{y}D(y_{t+1}\|y_{t})\rightarrow\infty$ if there exists $i\in[n]$ such that $y_{t+1,i}\rightarrow 0$.
@@
The detailed version of Theorem 8 and its proof can be found in Section E.2. Here since for all no-regret algorithms $\mathcal{A}_{2}$ we have $f(T)=\Omega(\sqrt{T})$, if we take $g(T)=\sqrt{Tf(T)}$, $\left(\frac{T}{g(T)}\right)^{2}=o(g(T))$ and we have $StackReg_{1}(\mathcal{A}_{1},\mathcal{A}_{2})=O(\sqrt{Tf(T)})$.
@@ 7 Conclusion
Acknowledgements
EM acknowledges support from NSF Award 2240110. YM is supported by: NSF Award CCF-2112665 (TILOS), DARPA AIE program, the U.S. Department of Energy, Office of Science, and CDC-RFA-FT-23-0069 from the CDC’s Center for Forecasting and Outbreak Analytics.
@@ B.1 Proof of Proposition 1
B.2 Refined Statement and Proof of Theorem 3
We first provide a refined statement of Theorem 3 that expands the big $O(\cdot)$ notation in the original statement.
Theorem 9.
If $P_{1}$ has a set of $d_{1}$-pessimistic facets $E_{i}^{-},\forall i\in[n]$ such that $\forall i\neq j$, $\inf_{x\in E_{i}^{-},x^{\prime}\in E_{j}}\|x-x^{\prime}\|_{1}\geq d_{2}$, as long as $P_{2}$ is using an $f$-no-regret algorithm $\mathcal{A}_{2}$ with constant $C$, $P_{1}$ can guarantee a Stackelberg regret of
$StackReg_{1}(\mathcal{A}_{1},\mathcal{A}_{2})=\left(Td_{1}+\frac{2Cf(T)}{\epsilon d_{2}}\right)\|A\|_{\max}$ (45)
by sticking to the corresponding $x^{-}$ obtained from (9) and (10). Here $\epsilon$ is a constant that depends only on the learner’s payoff matrix $B$.
Proof.
@@ B.2 Refined Statement and Proof of Theorem 3
where we have used the fact that $1-y_{t,i}=\sum_{j\in[n],j\neq i}y_{t,j}$ for all $y_{t}\in\Delta_{n}$ in the fourth equation. Since $\mathcal{A}_{2}$ used by the learner is $f$-no-regret, assume the regret constant is $C$, we have:
@@
$\leq$ $\frac{2Cf(T)}{\epsilon d_{2}},$
where the second equality follows from $y_{t,i}\in[0,1],\forall y_{t}\in\Delta_{n},i\in[n]$ and the third equality follows from the same argument as above. Let $i^{*}$ denote $\operatorname*{arg\,max}_{i}V_{i}(A,B)$, the Stackelberg regret of $P_{1}$ satisfies:
@@
$=$ $T\cdot\left(V_{i^{*}}(A,B)-V_{i^{*}}^{-}(A,B)\right)+\|A^{T}x^{-}\|_{\infty}\|\sum_{t=1}^{T}(e_{i^{*}}-y_{t})\|_{1}$
$\stackrel{{\scriptstyle\text{(iii)}}}{{\leq}}$ $Td_{1}\|A\|_{\max}+\|A^{T}x^{-}\|_{\infty}\|\sum_{t=1}^{T}(e_{i^{*}}-y_{t})\|_{1}$
$\stackrel{{\scriptstyle\text{(iv)}}}{{\leq}}$ $Td_{1}\|A\|_{\max}+\|A\|_{\max}\|\sum_{t=1}^{T}(e_{i^{*}}-y_{t})\|_{1}$
$\stackrel{{\scriptstyle\text{(v)}}}{{\leq}}$ $Td_{1}\|A\|_{\max}+2\|A\|_{\max}\frac{Cf(T)}{\epsilon d_{2}},$
where we have used Hölder’s inequality in (i), the maximizing argument of (10) in (ii), Proposition 1 in (iii), Hölder’s inequality in (iv) and (48) in (v). This completes the proof of Theorem 3. ∎
@@ C.4 Proof of Lemma 5
C.5 Refined Statement and Proof of Theorem 6
We first expand the big $O(\cdot)$ notation in the statement of Theorem 6 and obtain the following theorem:
Theorem 13.
Under 1, if $P_{1}$ has an estimator $\hat{\mathcal{B}}$ of $\mathcal{B}$ such that $\|d_{i}(\mathcal{B},\hat{\mathcal{B}})\|_{\infty}\leq\epsilon=O(g(T)/T),\forall i$ for some $g(T)=o(T)$, then if $P_{2}$ is using a $f$-no-regret algorithm $\mathcal{A}_{2}$ with constant $C$, there exists an algorithm $\mathcal{A}_{1}$ satisfying:
$StackReg_{1}(\mathcal{A}_{1},\mathcal{A}_{2})=$ $\left(4\left(\epsilon T+\sqrt{Tf(T)}\right)Sen\left((\mathcal{B}_{i^{*}}^{\circ})^{T}\right)+\frac{C\sqrt{Tf(T)}}{\max_{j,k}\|B(e_{j}-e_{k})\|_{\infty}}\right)\|A\|_{\max}$ (82)
$=$ $O\left(\sqrt{Tf(T)}+g(T)\right).$
Proof.
@@ C.5 Refined Statement and Proof of Theorem 6
$\text{maximize}_{i,x}$ $V^{-}_{i}(A,\hat{\mathcal{B}})=x^{T}A_{:,i}$ (83)
@@
$\leq$ $4\left(\epsilon T+\sqrt{Tf(T)}\right)\|A_{:,i^{*}}\|_{\infty}Sen\left((\mathcal{B}_{i^{*}}^{\circ})^{T}\right)+\|A\|_{\max}\frac{C\sqrt{Tf(T)}}{\max_{j,k}\|B(e_{j}-e_{k})\|_{\infty}}$
$\leq$ $\left(4\left(\epsilon T+\sqrt{Tf(T)}\right)Sen\left((\mathcal{B}_{i^{*}}^{\circ})^{T}\right)+\frac{C\sqrt{Tf(T)}}{\max_{j,k}\|B(e_{j}-e_{k})\|_{\infty}}\right)\|A\|_{\max}$
@@ Appendix E Algorithms and Proofs for Section 6
Let $\tilde{y}\in BR(B,\tilde{x})$ be a best response to $\tilde{x}$. For the “if” direction, on observing that:
@@
for some constant $C^{\prime}$ where we have used the fact that $\tau=O(f(T))$, and therefore $\{x_{t},y_{t}\}_{t=1}^{T}$ is $f$-no-regret. For the “only if” direction, notice that
@@
E.1 Algorithm 1, Detailed Version and Proof of Theorem 7 and Discussion
@@
We also provide the detailed version of Theorem 7 as follows:
Theorem 16.
Suppose $m=n=2$ and the payoff matrix $B$ does not contain identical columns. For some chosen parameter $d$, if either one facet is empty, or each facet has diameter at least $d$ and $P_{2}$ uses an ascent algorithm $\mathcal{A}_{2}$ that is $f$-no-regret, Algorithm 1 with accuracy margin $d$ achieves a Stackelberg regret of at most
$\left(4+\frac{C}{\epsilon_{1}}f(T)\right)\|A\|_{\max}$ (106)
if the learner has a strictly dominated action, where $\epsilon_{1}=\min_{x\in\Delta_{m}}x^{T}B(e_{1}-e_{2})$, and
$\left(-2\log d+6+2Td+\frac{2Cf(T)}{\epsilon_{2}d}\right)\|A\|_{\max}$ (107)
otherwise, where $\epsilon_{2}$ is a constant that depends only on $B$. Under either case, $C$ is a constant that depends only on the learner regret constant and payoff matrix. This indicates that the Stackelberg regret is at most $O(\frac{f(T)}{d}+dT-\log d)$ as long as $d=\Omega(\exp(-f(T)))$.
Proof.
@@
Now we analyze the Stackelberg regret against an $f$-no-regret algorithm $\mathcal{A}_{2}$. W.l.o.g suppose $e_{2}$ is the strictly dominated action, since $e_{2}$ is strictly dominated, we would have $x^{T}Be_{1}-x^{T}Be_{2}>0$ for all $x\in\Delta_{m}$, there exists a constant $\epsilon$ that depends only on $B$ (but not $d$), such that:
$\tilde{x}^{T}Be_{1}-\tilde{x}^{T}Be_{2}\geq\epsilon.$ (108)
@@
$\|\sum_{t=5}^{T}(e_{1}-y_{t})\|_{\infty}\leq\frac{C}{\epsilon}f(T).$ (110)
@@
$StackReg_{1}(\mathcal{A}_{1},\mathcal{A}_{2})$ $\leq 4V(A,B)-\sum_{t=1}^{4}x_{t}^{T}Ay_{t}+\tilde{x}^{T}A\sum_{t=5}^{T}(e_{1}-y_{t})$ (111)
$\leq 4\|A\|_{\max}+\|\tilde{x}^{T}A\|_{1}\|\sum_{t=5}^{T}(e_{1}-y_{t})\|_{\infty}$
$\leq 4\|A\|_{\max}+\|A\|_{\max}\frac{C}{\epsilon}f(T).$
If neither facets are empty, the binary search phase shrinks the interval length from $1$ to $d$, which requires at most $-\log d+1$ calls of the BinarySearch function. Each Binary search calls the procedure test for at most one time, which corresponds to at most two interaction time steps. Therefore, the total number of time steps in the binary search phase is at most $-2\log d+6$, leading to a Stackelberg regret of at most $(-2\log d+6)\|A\|_{\max}$, which is $O(f(T))$.
@@
$StackReg_{1}(\mathcal{A}_{1},\mathcal{A}_{2})$ $\leq\lfloor-2\log d+6\rfloor V(A,B)-\sum_{t=1}^{\lfloor-2\log d+6\rfloor}x_{t}^{T}Ay_{t}+\tilde{x}^{T}A\sum_{t=\lfloor-2\log d+6\rfloor+1}^{T}(e_{1}-y_{t})$ (112)
$\leq(-2\log d+6)\|A\|_{\max}+\|\tilde{x}^{T}A\|_{1}\|\sum_{t=\lfloor-2\log d+6\rfloor+1}^{T}(e_{1}-y_{t})\|_{\infty}$
$\leq(-2\log d+6)\|A\|_{\max}+\left(2Td+\frac{2Cf(T)}{\epsilon d}\right)\|A\|_{\max}.$
@@
E.2 Algorithm 4, Detailed Version and Proof of Theorem 8
@@
The detailed version of Theorem 8 is presented as follows:
Theorem 17.
If the learner payoff matrix $B$ statisfies the assumptions needed in Theorem 6, $P_{2}$ follows update rule (26), and each entry $\xi_{t,i}$ is i.i.d. $R$-sub-Gaussian, then with probability at least $1-\delta$, $P_{1}$ using Algorithm 4 with $k=\left(T/g(T)\right)^{2}2R^{2}\log(2mn/\delta)$, incurs Stackelberg regret of at most
$\left(m\left(1+k\right)+4\left(\frac{8ng(T)}{\max_{j,k}\|B(e_{j}-e_{k})\|_{\infty}}+\sqrt{Tf(T)}\right)Sen((\mathcal{B}_{i^{*}}^{\circ})^{T})+\frac{C\sqrt{Tf(T)}}{\max_{j,k}\|B(e_{j}-e_{k})\|_{\infty}}\right)\|A\|_{\max}.$ (118)
where $C$ is a constant that depends only on the learner regret constant and payoff matrix.
Proof.
@@
$L(y,\lambda,\mu)=\eta_{t}D(y\|y_{t})-(x_{t}^{T}B+\xi_{t}^{T})y-\lambda^{T}y+\mu(\sum_{i}y_{i}-1).$ (119)
Since $y_{t+1}$ is the optimal solution to (26), the KKT condition yields:
$\eta_{t}\nabla_{t+1}yD(y_{t+1}\|y_{t})-B^{T}x_{t}-\lambda+\mu\mathrm{1}-\xi_{t}=0;$ (120)
@@
Since the Bregman divergence regularizer satisfies $\nabla_{y}D(y_{t+1}\|y_{t})\rightarrow\infty$ if there exists $i\in[n]$ such that $y_{t+1,i}\rightarrow 0$, we can deduce that $\lambda=0$, and the condition becomes:
$\eta_{t}\nabla_{y}D(y_{t+1}\|y_{t})-B^{T}x_{t}+\mu\mathrm{1}-\xi_{t}=0.$ (121)
Let $h_{t}$ denote $h_{t}:=\eta_{t}\nabla_{y}D(y_{t+1}\|y_{t})$, we know that
@@
$StackReg_{1}(\mathcal{A}_{1},\mathcal{A}_{2})$ (132)
$\leq m(k+1)V(A,B)+\left(4\left(\frac{8ng(T)}{\max_{j,k}\|B(e_{j}-e_{k})\|_{\infty}}+\sqrt{Tf(T)}\right)Sen\left((\mathcal{B}_{i^{*}}^{\circ})^{T}\right)+\frac{C\sqrt{Tf(T)}}{\max_{j,k}\|B(e_{j}-e_{k})\|_{\infty}}\right)\|A\|_{\max}$
$\leq m(k+1)\|A\|_{\max}+\left(4\left(\frac{8ng(T)}{\max_{j,k}\|B(e_{j}-e_{k})\|_{\infty}}+\sqrt{Tf(T)}\right)Sen\left((\mathcal{B}_{i^{*}}^{\circ})^{T}\right)+\frac{C\sqrt{Tf(T)}}{\max_{j,k}\|B(e_{j}-e_{k})\|_{\infty}}\right)\|A\|_{\max}$
@@
Appendix F Numerical Experiments
F.1 Empirical Simulations for Section 6.1
For all experiments in this section, we assume the learner is using Online Gradient descent (OGD) with step size
$\eta_{t}=\frac{\eta_{0}}{\sqrt{t}}$ (133)
For the purpose of properly displaying the interaction and learning process, we choose different $\eta_{0}$ for different game instances. For each game instance, we compare the performance and learning dynamics for optimizer algorithm being either OGD or Binary Search explore-then-commit (BS, Algorithm 1). For Binary Search, we set the accuracy margin $d=0.01$. For each game instance, we plot both the payoff and the strategy (indicated by its 0-th entry) of each player at different time steps. We assume optimizer is the row player and learner is the column player.
Matching pennies.
We first test repeated matching pennies, where the payoff matrices are given by:
$A=\begin{bmatrix}1&-1\\ -1&1\end{bmatrix};B=\begin{bmatrix}-1&1\\ 1&-1\end{bmatrix}.$ (134)
Both the unique Nash equilibrium and the Stackelberg equilibria all have
$x=(\frac{1}{2},\frac{1}{2})^{T}.$ (135)
We obtain the curve shown in Figure 3.
Figure 3: Learning dynamics for optimizer algorithms OGD and BS for matching pennies.
Here the blue curves represent the learning dynamics when the optimizer uses binary search (Algorithm 1) with the exploration phase shaded in red. The orange curves represent the dynamics when the optimizer uses OGD. In both optimizer algorithms, the solid lines are curves for the optimizer and the dashed lines are curves for the learner. We plot the optimizer Stackelberg strategy and payoff in black dotted lines.
We can see that when both players are using OGD, the trajectory keeps oscillating and does not converge to the Nash equilibrium. In comparison, when the optimizer uses BS, it quickly learns its real underlying Stackelberg equilibrium (which is also the Nash) and commits to it, yielding a stable learning dynamics.
Constructed game instance 1.
Below we show that BS indeed yields a smaller Stackelberg regret than OGD. We construct the following game instance:
$A=\begin{bmatrix}5&0\\ 0&3\end{bmatrix};B=\begin{bmatrix}-2&2\\ 3&-3\end{bmatrix}.$ (136)
The unique Stackelberg equilibrium action for the optimizer is:
$x=(\frac{3}{5},\frac{2}{5})^{T}$ (137)
with Stackelberg value $3$. We obtain the curve shown in Figure 4.
Figure 4: Learning dynamics for optimizer algorithms OGD and BS for game instance 1.
In Figure 4 all curves are drawn with the same line style as in Figure 3, in addition we use blue and orange dotted lines to plot the average optimizer payoff for BS and OGD respectively.
We notice again that when the optimizer is using OGD, the algorithm fails to converge. In addition, after the optimizer commits to the pessimistic Stackelberg solution, the learner slowly converges to the best response induced by the Stackelberg equilibrium and steers the optimizer payoff close to the Stackelberg value, which is higher on average than the payoff using SGD.
Constructed game instance 2.
One may argue that OGD fails because it doesn’t converge, however it is not the case. Below we construct a game instance that has a unique Nash equilibrium to which OGD converges, and a unique Stackelberg equilibrium with higher optimizer utility than that of Nash. The game instance is as follows:
$A=\begin{bmatrix}2&0\\ 3&1\end{bmatrix};B=\begin{bmatrix}1&0\\ 0&2\end{bmatrix}.$ (138)
The unique Nash equilibrium is
$x=y=(0,1)^{T},$ (139)
while the unique Stackelberg equilibrium is
$x=(2/3,1/3)^{T},y=(1,0)^{T}.$ (140)
The optimizer payoff at Nash is $1$, while its Stackelberg value is $2$. The simulation result is shown in Figure 5.
Figure 5: Learning dynamics for optimizer algorithms OGD and BS for game instance 2.
The plot above shows that even if both converges, BS and OGD converge to completely different equilibria, while BS always yields a higher payoff and therefore a lower Stackelberg regret.
F.2 Empirical Simulations for Section 6.2
In this section we show the effectiveness of Algorithm 4. and illustrate the necessity of pessimism. Here we assume the optimizer is using Algorithm 4, but with different pessimism levels $d\in\{0.01,0.02,0.05\}$. We assume that the learner is using Stochastic Mirror descent with KL regularizer. For each pure strategy of the optimizer, we set the number of steps for exploration to be $k=50$. We consider the following game instance:
$A=\begin{bmatrix}0&1\\ 5&0\end{bmatrix};B=\begin{bmatrix}2&-2\\ -3&3\end{bmatrix}.$ (141)
The unique Stackelberg equilibrium of this game is
$x=(\frac{3}{5},\frac{2}{5})^{T}$ (142)
with optimizer payoff $2$. We plot the payoffs and strategies of both player at each time step with different $d$ in Figure 6.
Figure 6: Learning dynamics of payoff estimation for different pessimism levels $d$.
In Figure 5, different line colors indicate the learning dynamics of different pessimism levels. The solid lines are optimizer curves while the dashed lines being learner curves. The shaded region indicates the exploration phase and the black dotted line represents the optimizer Stackelberg payoff and strategy.
We can see that for larger $d$, the optimizer is being more pessimistic and chooses an action that is farther away from the Stackelberg equilibrium. This leads to a lower optimizer payoff after the learner converges to the unique best response induced by the action. However, for less pessimistic choices of $d$, since the committed optimizer strategy $\tilde{x}$ is too close to the Stackelberg equilibrium where the learner is indifferent from all mixed strategies, the gradients of the learner payoff with respect $\tilde{x}$ will be extremely small and thus takes a lot longer to converge. Once it hasn’t converged the optimizer payoff of committing to $\tilde{x}$ will be smaller, illustrating the effectiveness of being pessimistic.
Appendix G Proof of Auxiliary Lemmas
G.1 Proof of Lemma 10